\begin{references}





\bibitem{Sandiego} A.V.\ Balatsky, E.\ Abrahams, D.J.\
Scalapino and  J.R.\ Schrieffer, Proceedings of the Intl Conf. on
Strongly Correlated Electrons (SCEE), San Diego, 1993; Physica B, to
be published. See also cond-mat@babbage.sissa.it No 9309014.




\bibitem{Ber} V.L. Berezinskii, JETP Lett.\ {\bf 20}, 287 (1974);
[ZhETP Pisma {\bf 20}, 628 (1974)].



\bibitem{BA} A.V.\ Balatsky and  E.\ Abrahams, Phys.\ Rev.\ B {\bf
 45}, 13125 (1992).




\bibitem{G4} E.\ Abrahams, A.V.\ Balatsky, D.J.\ Scalapino and
J.R.\ Schrieffer, unpublished; also A. V. Balatsky and J. Bon\v ca,
Phys.\ Rev.\ B {\bf 48}, 7445 (1993).


\bibitem{EK} V.J.\ Emery and S.\ Kivelson, Phys.\ Rev.\ B{\bf 46},
10812 (1992).

\bibitem{C} P.\ Coleman {\it et.al.} unpublished, Phys. Rev. Lett. {\bf
72},  2960 (1993).



\end{references}

\documentstyle[preprint,revtex]{aps}

\begin{document}

\draft
\preprint{cond-mat/9312094}
\begin{title}
Erratum: Properties of Odd Gap Superconductors
\end{title}

\author{A. V. Balatsky}
\begin{instit}
 Theoretical Division\\ Los
Alamos National Laboratory, Los Alamos, NM 87545
\end{instit}
\author{Elihu Abrahams}
\begin{instit}
Serin Physics Laboratory, Rutgers University, P.O. Box
849, Piscataway, NJ 08855
\end{instit}
\author{D.J. Scalapino}
\begin{instit}
Department of Physics, University of California, Santa Barbara, CA
\end{instit}
\author {J.R. Schrieffer}
\begin{instit}
Department of Physics, Florida State University, Tallahassee, FL 32310
\end{instit}
\receipt{December 26 1993}
\begin{abstract}

Key words: superconductivity, odd frequency gap, composite operator.

\end{abstract}

\pacs{PACS Nos. 74.20-z;74.65+n}

In the recent paper \cite{Sandiego} an erroneous statement
that the composite order parameter for  odd-frequency pairing
was originally proposed by Berezinskii \cite{Ber} was
made.

Since the recent revival of the interest in
odd-frequency pairing \cite{BA},
the composite order parameter has been proposed in Abrahams et.al.
\cite{G4}, V. Emery and S. Kivelson \cite{EK} and in Coleman et.al.
\cite{C}.

\end{document}